# OpenMathInstruct-1: A 1.8 Million Math Instruction Tuning Dataset

### v1

@@ 4 Results
We finetune all the models on a mixture of (a) 512K fair downsampled GSM8K instances and (b) 512K MATH instances with any-code filtering (Section 2.4). [^10] The actual number of MATH instances is 511,677. Thus, the total finetuning corpus size is roughly 1.2M. We will justify the data selection choice later in the ablation experiments.
Table 3 compares the performance of OpenMath-finetuned models against their gpt-distilled counterparts. Among the 7B models, our OpenMath-Mistral-7B is competitive with all the gpt-distilled models. It is second-best to WizardMath on GSM8K, and bested by ToRA by 0.1% on MATH. [^11] Our grading script scores the publicly released ToRA outputs about 2-3% lower than the reported numbers. We believe that ToRA uses some heuristics to extract answers when the model doesn’t generate answers in the correct format. Our models easily outperform both MAmmoTH and MetaMath, even when controlling for the base fine-tuned model. Since WizardMath and ToRA finetuning datasets are not publicly available yet, OpenMathInstruct-1 presents a superior alternative to the publicly available MetaMathQA and MathInstruct datasets, which are used to fine-tune MetaMath and MAmmoTH, respectively.
With the increase in model parameters, our models continue to outperform MAmmoTH and MetaMath substantially. Compared to ToRA, with greedy decoding, we see a meaningful drop in performance on MATH, though our models are equal or better on GSM8K. With self-consistency (SC) decoding, however, our models outperform ToRA on both MATH and GSM8K. The substantial gains with SC can be attributed to the diversity of our fine-tuning data.
@@ 4.1 Ablations
Table 4: Comparison of performance of fair vs naive sampling on our validation subset of GSM8K and MATH.
@@
Random 74.3 35.0
Fair 75.3 37.0
4.1.1 Fair vs Naive Downsampling
@@ 4.1.2 Impact of Fine-Tuning Dataset Size
To determine the impact of the size of the fine-tuning dataset, we create datasets of size 128K/256K/512K by combining 64K/128K/256K fair downsampled subsets of GSM8K and MATH. Table 5 shows that the performance increases on both GSM8K and MATH with the increase in the fine-tuning dataset size. We didn’t find benefit from training the models for more steps, so the performance gain is attributable to the increased data size.
@@ 4.1.3 MATH-only Ablations
This section presents the ablation results for only the MATH portion of OpenMathInstruct-1. In all experiments, we finetune the base model on a 128K fair downsampled subset to control for the impact of data size.
@@
Default vs Subject-Specific Prompting.
In section 2.2.2, we motivated using subject-specific prompts, which ultimately didn’t result in much training set coverage difference. But how are the solutions generated by the combination of subject-wise prompts different from a single default prompt? To answer this, we create a subset of 128K instances generated with the default prompt/subject-specific prompts.
Table 6 compares the finetuning performance on these two splits on our MATH validation subset. While the model trained on the subject-specific subset trails the model trained on the default subset for greedy decoding; the trend is decisively reversed for self-consistent decoding with four samples. This suggests that the subset collected with subject-specific prompts has a higher diversity of solutions than the ones collected using a single prompt.
Code-Preferential Subsets.
@@
In this ablation, we determine the impact of code-preferential solution selection strategies proposed in Section 2.4.2. Table 7 shows that code-preferential solution strategies aid the greedy decoding performance. However, the reduction in solution diversity arguably results in decreased performance with self-consistency decoding (text-based solutions are only 1/3rd of the original corpus to begin with). Based on these results and because any-code results in a smaller finetuning dataset (512K compared to 664K with majority-code), we chose to use the any-code subset in our finetuning data blend.
@@ 5 Analysis
(a) Subject-wise performance
(b) Level-wise performance
Figure 5: Performance split by subjects and levels on our MATH validation subset.
@@
We analyze the performance of the ablation model trained on 512K instances from Section 4.1.2. We focus our discussion on the MATH benchmark where this model scores 41.6% on our MATH validation subset.
5.1 Performance-split by Subjects and Levels
Figure 5 presents the performance split by subjects and levels on the MATH validation subset. Among subjects, we see that the model’s worst performance is on geometry, which can be attributed to the lack of multi-modality in our base models Zhou et al. (2024). We see a monotonic decrease in performance with the increase in hardness level which is to be expected Zhou et al. (2024). The model scores 72.4% on Level 1 problems and only 16.3% on the hardest problems, i.e., Level 5.
5.2 Error Analysis
@@
Code Execution Error 78
Code timeout 15
Max code executions reached 10
@@
Table 9 presents the count of different error categories. For code-based solutions, we find that almost 74% of the errors in such solutions are due to reasoning error and the remaining 26% attributable to execution-related issues. We present sample solutions from these error categories in Appendix B.3.
@@ 6 Related Work
Recently, a plethora of work has been done on enhancing the mathematical reasoning capabilities of LLMs. Inference techniques such as Chain-of-Thought (Wei et al., 2022), its programmatic counterpart, Program of Thought (Gao et al., 2023; Chen et al., 2023b), Self-Consistency (Wang et al., 2023), and Self-Verification (Zhou et al., 2024) have been shown to significantly improve the reasoning capabilities of LLMs without any further training.
Pretraining language models on math-heavy content has resulted in foundation LLMs such as Minerva (Lewkowycz et al., 2022), Galactica (Taylor et al., 2022), and Llemma (Azerbayev et al., 2023) with stronger mathematical skills out-of-the-box. A more direct approach of dataset-specific training does instruction fine-tuning on problem-solution pairs derived from math reasoning datasets. Our work falls in this latter category and bears similarity with recent work such as RFT (Yuan et al., 2023), ToRA (Gou et al., 2024), MAmmoTH (Yue et al., 2024), MetaMath (Yu et al., 2024) and MathCoder (Wang et al., 2024). We differ from the previous work along one factor or a combination of the following factors: (a) reliance on GPT-3.5/4, (b) solution format, and (c) use of ground truth text solution in synthesizing code-based solutions.
@@
Recent work exploring the use of targeted synthetic data generated by large foundation models for pre-training/instruction tuning smaller LLMs has led to tremendous progress in reasoning skills of these smaller LLMs (Gunasekar et al., 2023; Li et al., 2023; Eldan and Li, 2023; Mukherjee et al., 2023; Xu et al., 2023; Liu et al., 2023).
@@ 7 Conclusion
We introduce OpenMathInstruct-1, a math instruction tuning dataset with 1.8M problem-solution pairs with a commercially permissive license. Compared to previous work, OpenMathInstruct-1 is at least four times bigger. The problems are taken from the training set of GSM8K and MATH benchmarks, and the solutions are synthesized by few-shot prompting the Mixtral model. With our proposed prompting novelty of using masked text solutions and some brute-force scaling, we achieve training set coverage of 99.9% for the GSM8K benchmark and 93% for the challenging MATH benchmark. The quality of these synthesized solutions is illustrated by finetuning experiments, which show models achieving performance comparable to or outperforming their gpt-distilled counterparts. To support the open-source efforts in this direction, we publicly release all our fine-tuned models, code, and the OpenMathInstruct-1 along with a further 6.6M incorrect sampled solutions.
Acknowledgement
We want to thank the NeMo team at NVIDIA for their support.

### v2

@@ 4 Results
We finetune all the models on a mixture of (a) 512K fair downsampled GSM8K instances, and (b) 512K MATH instances with any-code filtering (Section 2.4). [^10] The actual number of MATH instances is 511,677. Thus, the total finetuning corpus size is roughly 1.02M. We justify the data selection choices later in the ablation experiments.
Table 3 compares the performance of OpenMath-finetuned models against their GPT-distilled counterparts. Among the 7B models, our OpenMath-Mistral-7B is competitive with all the GPT-distilled models. It is second-best to WizardMath on GSM8K, and bested by ToRA by 0.1% on MATH. [^11] Our grading script scores the publicly released ToRA outputs about 2-3% lower than the reported numbers. We believe that ToRA uses some heuristics to extract answers when the model doesn’t generate answers in the correct format. Our models easily outperform both MetaMath [1] and MAmmoTH [2], even when controlling for the base fine-tuned model. Since WizardMath and ToRA finetuning datasets are not publicly available yet, OpenMathInstruct-1 presents a superior alternative to the publicly available MetaMathQA and MathInstruct datasets, which are used to fine-tune MetaMath and MAmmoTH, respectively.
With the increase in model parameters, our models continue to outperform MetaMath and MAmmoTH substantially. Compared to ToRA, with greedy decoding, we see a meaningful drop in performance on MATH, though our models are equal or better on GSM8K. With self-consistency (SC) decoding, however, our models outperform ToRA on both MATH and GSM8K. The substantial gains with SC can be attributed to the diversity of our fine-tuning data.
@@ 4.1 Ablations
Table 4: Comparison of performance of Fair vs Naive downsampling on our validation subset of GSM8K and MATH.
Sampling GSM8K MATH
Naive 74.3 35.0
Fair 75.3 37.0
Table 5: Comparison of Default vs Masked prompting on our validation subset of GSM8K and MATH.
@@
Default 73.8 36.9
Masked 77.7 37.4
4.1.1 Fair vs. Naive Downsampling
@@
4.1.2 Default vs Masked Prompting
We finetune the base model on a dataset of 128K instances created by combining 64K fair-sampled instances from the GSM8K and MATH portion of the data generated using default and masked prompting. Table 5 shows that the model fine-tuned on the data generated using masked prompting outperforms the one created by default prompting on both GSM8K and MATH. Thus, the gains in the training set coverage with masked prompting (Section 2.2.3) also translate to finetuning performance.
@@ 4.1.3 Impact of Fine-Tuning Dataset Size
To determine the impact of the size of the fine-tuning dataset, we create datasets of size 128K/256K/512K by combining 64K/128K/256K fair downsampled equally-sized subsets of GSM8K and MATH. Table 6 shows that the performance increases for both GSM8K and MATH with the increase in the fine-tuning dataset size. We didn’t find any benefit from training the models for more steps, so the performance gain is attributable to the increased data size.
@@ 4.1.4 MATH-only Ablations
This section presents the ablation results for only the MATH portion of OpenMathInstruct-1. We finetune the base model on a 128K fair downsampled subset to control for data size.
@@
Default vs Subject-Specific Prompting.
In section 2.2.2, we motivated using subject-specific prompts, which ultimately didn’t result in much training set coverage difference. But how are the solutions generated by the combination of subject-wise prompts different from a single default prompt? To answer this, we create a subset of 128K instances generated with the default prompt/subject-specific prompts. Table 7 compares the finetuning performance on these two splits on our MATH validation subset. While the model trained on the subject-specific subset underperforms the model trained on the default subset with greedy decoding, the trend is decisively reversed for self-consistency decoding with four samples. This suggests that the subset collected with subject-specific prompts has a higher diversity of solutions than the ones collected using a single prompt.
Code-Preferential Subsets.
In this ablation, we determine the impact of code-preferential solution selection strategies proposed in Section 2.4.2. Table 8 shows that code-preferential solution strategies aid the greedy decoding performance. However, the reduction in solution diversity arguably results in a performance drop with self-consistency decoding (text-based solutions are about one-third of the original corpus). Based on these results and because Any-Code results in a smaller finetuning dataset (512K compared to 664K with Majority-Code), we chose to use the Any-Code subset.
@@ 5 Analysis
We analyze the performance of the ablation model trained on 512K instances from Section 4.1.3. We limit the discussion to the MATH benchmark, where the model scores 41.6% on our validation subset.
Performance-split by Subjects and Levels.
Figure LABEL:fig:math_subj_level presents the performance split by subjects and levels on the MATH validation subset. Among subjects, we see that the model’s worst performance is on geometry, which can be attributed to the lack of multi-modality in our base models [23]. We see a monotonic decrease in performance with the increase in hardness level which is to be expected [23]. The model scores 72.4% on Level 1 problems and only 16.3% on the hardest problems, i.e., Level 5.
Error Analysis.
@@
Code Execution Error 178
Code timeout 115
Max code executions reached 110
@@
Table 10 presents the count of different error categories. For code-based solutions, we find that almost 74% of the errors in such solutions are due to reasoning errors, and the remaining 26% are attributable to execution-related issues. We present sample solutions from these error categories in Appendix B.3.
@@ 6 Related Work
Recently, a plethora of work has been done on enhancing the mathematical reasoning capabilities of LLMs. Inference techniques such as Chain-of-Thought [33], its programmatic counterpart, Program of Thought [24, 34], Self-Consistency [32], and Self-Verification [23] have been shown to significantly improve the reasoning capabilities of LLMs.
Pretraining language models on math-heavy content has resulted in foundation LLMs such as Minerva [35], Galactica [36], Llemma [37], and DeepSeekMath [38] with stronger mathematical skills out-of-the-box. A more direct approach of dataset-specific training does instruction fine-tuning on problem-solution pairs derived from math reasoning datasets. Our work falls in this latter category and bears similarity with recent work such as RFT [39], ToRA [8], MAmmoTH [2], MetaMath [1] and MathCoder [7]. We differ from the previous work along one factor or a combination of the following factors: (a) reliance on GPT-3.5/4, (b) solution format, and (c) use of ground truth text solution in synthesizing code-based solutions.
@@
Recent work exploring the use of targeted synthetic data generated by large foundation models for pre-training/instruction tuning smaller LLMs has led to tremendous progress in skills of these smaller LLMs [5, 6, 4, 40, 41, 42].
@@ 7 Conclusion
We introduce OpenMathInstruct-1, a math instruction tuning dataset with 1.8M problem-solution pairs which is released with a commercially permissive license. Compared to previous work, the OpenMathInstruct-1 dataset is at least four times bigger. With our proposed prompting novelty of using masked text solutions and some brute-force scaling, we achieve training set coverage of 99.9% for the GSM8K benchmark and 93% for the challenging MATH benchmark. The quality of these synthesized solutions is illustrated by finetuning experiments, which show models achieving performance comparable to or better than their gpt-distilled counterparts. To support the open-source efforts in this direction, we publicly release all our fine-tuned models, code, and the OpenMathInstruct-1 along with a further 6.6M incorrect sampled solutions.
Limitations and Potential Risks
Our work aims to improve the mathematical reasoning of open-source models using open-source models. In pursuit of this goal, we create a synthetic dataset, OpenMathInstruct-1, that our experiments show aids the performance on existing math benchmarks. Below, we list the key limitations of our work:
  • Our manual analysis reveals solutions that get the right answer but via flawed reasoning (Figure 14 in Appendix). Removing these semantically noisy solutions is beyond the scope of the current work. This means a lack of guarantee about the quality of our synthetically generated solutions.
  • Improving performance on in-domain math benchmarks may not translate to performance gain on other related tasks. The drop in performance on GSM-Hard compared to GSM indicates that our models may not be robust to input perturbations, though, they are at par with previous work.
While we don’t foresee any material risk due to our work, using our imperfect dataset and models to perform tasks, such as evaluating student assignments or building a math tutor, carries risk.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our claims match our experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  10. 3.
    Theory Assumptions and Proofs
  11. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  12. Answer: [N/A]
  13. Justification: Empirical paper
  14. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  15. 4.
    Experimental Result Reproducibility
  16. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  17. Answer: [Yes]
  18. Justification: We provide all the details to the best of our capacity. Some of the details have been shared in the supplementary material.
  19. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  20. 5.
    Open access to data and code
  21. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  22. Answer: [No]
  23. Justification: We are not sharing the data and code as of now. But we will be releasing everything with open access after the review cycle.
  24. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  25. 6.
    Experimental Setting/Details
  26. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  27. Answer: [Yes]
  28. Justification: All the details have been shared in the paper.
  29. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  30. 7.
    Experiment Statistical Significance
  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  32. Answer: [No]
  33. Justification: All our fine-tuning runs are quite costly, which limits our capability to do multiple runs to capture this variability.
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  35. 8.
    Experiments Compute Resources
  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  37. Answer: [Yes]
  38. Justification: Details in the Appendix.
  39. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  40. 9.
    Code Of Ethics
  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  42. Answer: [Yes]
  43. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  44. 10.
    Broader Impacts
  45. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  46. Answer: [Yes]
  47. Justification: We briefly mention it in the Limitations section on Page 10.
  48. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  49. 11.
    Safeguards
  50. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  51. Answer: [No]
  52. Justification: Our dataset is related to mathematical reasoning.
  53. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  54. 12.
    Licenses for existing assets
  55. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  56. Answer: [Yes]
  57. Justification: We mention the use of Mixtral models and that these models are released under permissive license.
  58. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  59. 13.
    New Assets
  60. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  61. Answer: [Yes]
  62. Justification: We have tried to share details of the dataset generation process and the dataset ultimately created to the best of our capabilities. We have also clearly communicated that all the resources will be released under a commercially permissive license.
  63. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  64. 14.
    Crowdsourcing and Research with Human Subjects
  65. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  66. Answer: [N/A]
  67. Justification: Not a paper which deals with human subjects.
  68. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  69. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  70. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  71. Answer: [N/A]
  72. Justification: Not required for our paper.
  73. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.